# Mitigating Graph Covariate Shift via Score-based Out-of-distribution Augmentation

### v1

@@ Abstract
Distribution shifts between training and testing datasets significantly impair the model performance on graph learning. A commonly-taken causal view in graph invariant learning suggests that stable predictive features of graphs are causally associated with labels, whereas varying environmental features lead to distribution shifts. In particular, covariate shifts caused by unseen environments in test graphs underscore the critical need for out-of-distribution (OOD) generalization. Existing graph augmentation methods designed to address the covariate shift often disentangle the stable and environmental features in the input space, and selectively perturb or mixup the environmental features. However, such perturbation-based methods heavily rely on an accurate separation of stable and environmental features, and their exploration ability is confined to existing environmental features in the training distribution. To overcome these limitations, we introduce a novel approach using score-based graph generation strategies that synthesize unseen environmental features while preserving the validity and stable features of overall graph patterns. Our comprehensive empirical evaluations demonstrate the enhanced effectiveness of our method in improving graph OOD generalization.
1 Introduction
Deep learning algorithms have become predominant in the analysis of graph-structured data. However, a common limitation of existing methods is the assumption that both training and testing graphs are independently and identically distributed (i.i.d.). This assumption often falls short in real-world scenarios, where shifts in data distribution frequently occur, leading to significant degradation in model performance. As a result, there has been considerable progress in improving graph out-of-distribution (OOD) generalization, evidenced by advancements in invariant graph learning (Chen et al., 2022; Wu et al., 2022; Huang et al., 2024) and graph data augmentation (Rong et al., 2019; Wang et al., 2021; Han et al., 2022; Yao et al., 2022; Sui et al., 2024; Li et al., 2024).
Recent studies (Gui et al., 2022; Sui et al., 2024) have identified two primary types of distribution shifts. Correlation shifts occur when the statistical relationships between environments and labels differ between the training and testing datasets, assuming that the test environments are represented within the training dataset. Covariate shifts, on the other hand, arise when the test environments are not present in the training dataset. A prevalent causal perspective in graph invariant learning  (Chen et al., 2022; Wu et al., 2022) suggests that stable features of graphs, which causally determine labels, remain invariant across different environments, whereas varying environmental features contribute to distribution shifts. Consequently, previous studies (Miao et al., 2022; Chen et al., 2022; Wu et al., 2022) have primarily focused on correlation shifts by isolating invariant stable graph patterns from the environmental features. In this work, we focus on the relatively neglected but challenging problem of covariate shift in graph learning.
While different graph data augmentation techniques have been proposed to generate new data across various domains (Rong et al., 2019; Han et al., 2022; Sui et al., 2024), these methods mainly modify existing data within the training set by mixing or dropping edges, which either have limited environmental exploration ability or could result in invalid data samples (e.g., molecular graphs violating chemical rules). Moreover, indiscriminate augmentations (Rong et al., 2019) can distort stable patterns, resulting in uncontrollable augmented distributions. While controlled augmentations (Sui et al., 2024) have shown promising outcomes, they heavily depend on the accurate separation of stable and environmental patterns, which remains a nontrivial challenge and could be inherently infeasible. These observations prompt an essential inquiry: ”Is it possible to explore the training distribution under control, such that the exploration extends beyond the confines of the training environments while still preserving essential stable patterns?
In response, this work introduces an innovative score-based graph augmentation strategy that mitigates graph covariate shift by improving the exploration of training distribution while preserving stable predictive patterns on the generated graphs. In high level, we formulate the problem of OOD data augmentation as graph generation simultaneously conditioned on graph labels and exploration variables, based on the graph generation hypothesis widely used in prior studies (Wu et al., 2022; Yang et al., 2022; Gui et al., 2022; Chen et al., 2022). Specifically, we employ a score-based diffusion probabilistic model, commonly known as a diffusion model (Song et al., 2020), to effectively capture the data distribution of unlabeled graphs. During the generation phase, we introduce a novel guidance scheme that generates augmented graphs, concurrently retaining predictive stable patterns and incorporating explored environments. Our proposed Out-of-Distribution Diffusion Augmentation (OODA) framework utilizes graph labels to guide the sampling process toward graphs that are highly likely to contain stable patterns. The exploration parameter facilitates exploration beyond the training graph space by flexibly adjusting the discrepancy from the training distribution. The robustness of the score-based diffusion model ensures the validity of the generated graphs, preventing the formation of invalid structures (e.g., non-viable molecules) that could impair downstream classification performance. Furthermore, our guidance scheme eliminates the need to explicitly split graphs into stable and environmental subgraphs. We experimentally validate our method on both synthetic and real-world graph classification tasks under diverse covariate shift settings. Our results demonstrate that OODA outperforms state-of-the-art baselines, including invariant graph learning and graph data augmentation, highlighting its effectiveness in exploring environments under control while preserving stable patterns.
@@ 1 Introduction
  • We propose a novel graph generation-based environment augmentation approach to address covariate distribution shifts in graph learning. Our method enables controlled exploration of environmental patterns while preserving stable patterns, without the need to explicitly separate them.
  • Our approach can simultaneously generate out-of-distribution (OOD) graph structures, node features, and edge features, making it uniquely capable of handling covariate shifts in both feature and structural distributions, as well as when these shifts occur simultaneously.
  • Extensive empirical evaluations demonstrate that our framework outperforms state-of-the-art graph OOD generalization methods across diverse tasks, including synthetic, semi-artificial, real-world molecular, and natural language sentiment analysis datasets.
2 Related Work
@@
Graph-structured data are inherently complex, characterized by the intricate challenges of irregularity and nuanced structural information. This complexity gives rise to graph out-of-distribution (OOD) problems that not only necessitate addressing shifts in feature distributions but also demand attention to variations of structural distributions. In this context, we summarize two principal categories of algorithms for graph OOD robustness: (i) invariant graph learning strategies, which aim to ensure model stability across varying distributions; and (ii) graph data augmentation techniques, designed to enhance model generalizability by simulating diverse distribution scenarios.
@@
Invariant Graph Learning. The concept of invariant graph learning draws inspiration from seminal works such as those by (Arjovsky et al., 2019; Rosenfeld et al., 2020; Ahuja et al., 2021). This approach aims at identifying stable graph structures (e.g., subgraphs) or representations (predictors) that remain consistent across different environments, thereby enhancing out-of-distribution (OOD) generalization. This is achieved by capturing salient graph features and minimizing empirical risks across varying conditions. For example, DIR (Wu et al., 2022) distinguishes between invariant and environment-specific subgraphs by creating varied interventional distributions on the training distribution. CIGA (Chen et al., 2022) further explores this domain by employing synthetic environments and the graph generation process to identify stable features under various distribution shifts. However, this line of research assumes access to test environments during training, which is an unrealistic assumption given the impracticality of covering all possible test scenarios. Training in limited environments reduces spurious correlations but fails to generalize to new, unseen environments. DISGEN (Huang et al., 2024) gains promising results in disentangling the size factors from graph representations by minimizing the shared information between size- and task-related information, however, the technique is constrained to handle size generalization. In this work, we propose a framework capable of generalizing to unseen environments characterized by differences not only in graph size but also in graph structure, node features, and edge features.
@@
Graph Data Augmentation. Beyond invariant graph learning, graph data augmentation aims to diversify the training distribution, thereby enhancing the out-of-distribution (OOD) generalization of models. DropEdge (Rong et al., 2019) introduces randomness by selectively removing edges, thus varying the training data’s structure. M-Mixup (Wang et al., 2021) enriches the dataset by interpolating diverse and irregular graphs within semantic space. $\mathcal{G}$-Mixup (Han et al., 2022) extends this concept to graph classification, interpolating across different graph generators (graphons) to produce augmented graphs. Adversarial augmentation techniques, such as FLAG (Kong et al., 2022), apply gradient-based perturbations to node features, and AIA (Sui et al., 2024) generates adversarial masks on graphs, both aimed at probing environmental discrepancies. Despite these advancements, overcoming the limitations in environmental exploration caused by modifying graphs within the original training set remains a challenge, indicating ongoing opportunities for innovation in graph data augmentation strategies. Recently, environment-aware augmentation frameworks (Li et al., 2024) have utilized environment information to linearly explore training graph structures and node features, however, they depend heavily on high-quality and sufficiently diverse environment information. In practice, annotating environment labels or capturing diverse environment information is costly and often infeasible. In this work, we introduce a generation-based augmentation method that eliminates the need for accessing environment information.
3 Problem Formulation
@@
Notations. We represent a graph with $n$ nodes as $G=(\bm{A},\bm{X},\bm{E})$, where $\bm{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, $\bm{X}\in\mathbb{R}^{n\times a}$ denotes $a$-dimensional node features and $\bm{E}\in\mathbb{R}^{n\times n\times b}$ encodes $b$-dimensional edge features. Without the loss of generality, we focus on the graph classification task where each graph $G$ is associated with a label $Y\in\mathcal{Y}$, determined by a predefined labelling rule $\mathcal{G}\rightarrow\mathcal{Y}$. Following invariant learning (Ahuja et al., 2021; Chen et al., 2022), we denote the graph dataset as $\mathcal{D}=\{(G_{i}^{e},Y_{i}^{e})\}_{e\in\mathcal{E}_{\mathrm{all}}}$, where $(G_{i}^{e},Y_{i}^{e})\sim P_{e}(G,Y)$ is an i.i.d. draw in the environment $e$ sampled from all possible environments $\mathcal{E}_{\mathrm{all}}$. The complete dataset can be partitioned into a training set $\mathcal{D}_{\mathrm{tr}}=\{(G_{i}^{e},Y_{i}^{e})\}_{e\in\mathcal{E}_{\mathrm{tr}}}$ and a test set $\mathcal{D}_{\mathrm{te}}=\{(G_{i}^{e},Y_{i}^{e})\}_{e\in\mathcal{E}_{\mathrm{te}}}$, where $\mathcal{E}_{\mathrm{tr}}$ and $\mathcal{E}_{\mathrm{te}}$ index the training and testing environments, respectively. In practice, environment information may not be explicitly given, and we further denote the training distribution as $P_{\mathrm{tr}}(G,Y)$ and the testing distribution as $P_{\mathrm{te}}(G,Y)$.
@@
Graph Classification under Covariate Shift. With only observing the training set $\mathcal{D}_{\mathrm{tr}}$ sampled from the training distribution $P_{\mathrm{tr}}$ in training environments $\mathcal{E}_{\mathrm{tr}}$, our generalization objective under graph covariate shift is to train an optimal graph classifier $f:\mathcal{G}\rightarrow\mathcal{Y}$ that performs well across any possible environments $\mathcal{E}_{\mathrm{all}}\supseteq\mathcal{E}_{\mathrm{tr}}$. We formulate this goal as the following minimization problem:

### v2

# Score-based Conditional Out-of-Distribution Augmentation for Graph Covariate Shift

@@ Abstract.
Distribution shifts between training and testing datasets significantly impair the model performance on graph learning. A commonly-taken causal view in graph invariant learning suggests that stable predictive features of graphs are causally associated with labels, whereas varying environmental features lead to distribution shifts. In particular, covariate shifts caused by unseen environments in test graphs underscore the critical need for out-of-distribution (OOD) generalization. Existing graph augmentation methods designed to address the covariate shift often disentangle the stable and environmental features in the input space, and selectively perturb or mixup the environmental features. However, such perturbation-based methods heavily rely on an accurate separation of stable and environmental features, and their exploration ability is confined to existing environmental features in the training distribution. To overcome these limitations, we introduce a novel distributional augmentation approach enabled by a tailored score-based conditional graph generation strategies to explore and synthesize unseen environments while preserving the validity and stable features of overall graph patterns. Our comprehensive empirical evaluations demonstrate the enhanced effectiveness of our method in improving graph OOD generalization.
$c$s: Computing methodologies Neural networksccs: Mathematics of computing Graph algorithms
1. Introduction
Deep learning has become the dominant paradigm for analyzing graph-structured data across diverse domains. However, most graph learning methods assume that training and test graphs are independently and identically distributed (i.i.d.)., an assumption that frequently fail in real-world setting (gui2022good; huang2024enhancing). On the Web, spanning hyperlink graphs, social networks, recommendation and e-commerce interaction graphs, and knowledge graphs, distribution shift frequently arise due to product launches, policy or UI changes, seasonality, bot activity, and rapid community growth. Prior studies have shown that such shifts can lead to substantial performance degradation (gui2022good; sui2024unleashing), motivating growing interest in graph out-of-distribution (OOD) generalization, including invariant graph learning (chen2022learning; wu2022discovering; huang2024enhancing) and graph data augmentation (rong2019dropedge; wang2021mixup; han2022g; yao2022improving; sui2024unleashing; ligraph).
Following (gui2022good; sui2024unleashing), we distinguish correlation shift, where environment–label correlations differ but test environments are represented during training, from covariate shift, where the test environments are not present in the training dataset. A causal view in invariant graph learning (chen2022learning; wu2022discovering) posits that label-determining stable patterns are invariant across environments, while environmental patterns vary. Prior work largely targets correlation shift by isolating stable patterns (miao2022interpretable; chen2022learning; wu2022discovering). We focus instead on the underexplored covariate shift in graph learning, which is pervasive in Web settings (e.g., emerging communities or new interaction regimes unseen at train time).
Existing graph augmentation techniques typically perturb graph structures or features by mixing or dropping edges (rong2019dropedge; han2022g; sui2024unleashing). While effective with regularized contexts, such operations offer limited exploration ability to unseen environments and risk generating semantically inconsistent or unreliable samples (e.g., invalid molecules in chemistry), as indiscriminate edits may distort label-relevant stable patterns, yielding uncontrolled augmented distributions. Controlled methods (sui2024unleashing) help but often require an explicit, accurate split of stable vs. environmental subgraphs—nontrivial and sometimes infeasible on Web graphs. This raises a key question:
Can we explore beyond the training environments in a controllable way while preserving label-critical stable patterns?
We provide an affirmative answer via a score-based conditional graph generation strategy for covariate-shift mitigation. We formulate OOD augmentation as generating graphs conditioned on a target label and an exploration parameter that controls the degree of distributional deviation, building on the widely used graph-generation hypothesis (wu2022discovering; yang2022learning; gui2022good; chen2022learning). Inspired by diffusion models (song2020score), we define a forward corruption process and a novel reverse denoising process that jointly (i) retain predictive stable patterns and (ii) inject environment exploration. Our conditional score function incorporates both label information and the degree of OOD-ness, enabling Conditional Out-of-Distribution Diffusion Augmentation (CODA) to generate graphs likely to preserve stable patterns while probing underrepresented (low-density) regions [^1] Low-density regions are parts of the class-conditional data manifold with few samples; see Section 4.1. of the class-conditioned manifold beyond the training graph space.
CODA inherits diffusion-model robustness, helping ensure validity (preventing, e.g., non-viable molecules) without explicitly separating stable from environmental subgraphs. We evaluate on synthetic and real-world graph classification tasks under diverse covariate shifts, which are commonly observed in Web contexts. For instance, motif-frequency variations emulate evolving community or interaction structures; changes in graph size correspond to thread-depth or document-length drift; and feature/structural shifts reflect the dynamics of user/item attributes and connectivity patterns in online platforms. Across these settings, CODA achieves consistent gains over state-of-the-art invariant-learning and augmentation methods, enabling controlled exploration of OOD environments while preserving stable patterns.
@@ 1. Introduction
  • We propose a novel conditional graph generation-based augmentation framework that addresses covariate distribution shifts. Our method enables controlled exploration of environmental patterns while preserving stable patterns, without requiring explicit subgraph decomposition.
  • We develop a unified generator for OOD structures, node features, and edge features, handling feature, structural, and joint shifts which are pervasive in Web contexts.
  • Extensive experiments across synthetic, semi-artificial, molecular, and Web-analogous text/review graph datasets demonstrate consistent gains over strong OOD generalization baselines.
2. Related Work
@@
Graph-structured data, such as hyperlink graphs, social networks and e-commerce interaction graphs in Web, are inherently complex, characterized by the intricate challenges of irregularity and nuanced structural information. This complexity gives rise to graph OOD problems that not only necessitate addressing shifts in feature distributions but also demand attention to variations of structural distributions. In this context, we summarize two principal categories of algorithms for graph OOD robustness: (i) invariant graph learning strategies, which aim to ensure model stability across varying distributions; and (ii) graph data augmentation techniques, designed to enhance model generalizability by simulating diverse distribution scenarios.
@@
Invariant Graph Learning. The concept of invariant graph learning, inspired from seminal works (arjovsky2019invariant; rosenfeld2020risks; ahuja2021invariance), targets on identifying stable graph structures (e.g., subgraphs) or representations (predictors) that remain consistent across different environments, thereby enhancing OOD generalization. This is achieved by capturing salient graph features and minimizing empirical risks across varying conditions. In scenarios where establishing causality is complex or where strong assumptions may not hold, the task can be approximated by identifying features that demonstrate invariance under distributional shifts, thereby facilitating OOD generalization (li2022out). Effective OOD generalization is achieved by basing predictions solely on invariant information (li2022out). For example, DIR (wu2022discovering) distinguishes between invariant and environment-specific subgraphs by creating varied interventional distributions on the training distribution. CIGA (chen2022learning) further explores this domain by employing synthetic environments and the graph generation process to identify stable features under various distribution shifts. However, this line of research assumes access to test environments during training, which is an unrealistic assumption given the impracticality of covering all possible test scenarios. Training in limited environments reduces spurious correlations but fails to generalize to new, unseen environments. DISGEN (huang2024enhancing) gains promising results in disentangling the size factors from graph representations by minimizing the shared information between size- and task-related information, however, the technique is constrained to handle size generalization. Recently, CANET (wu2024graph) introduces an environment estimator that infers a pseudo-environment, enabling the model to capture stable features under node-level distribution shifts without requiring prior knowledge of environment labels. Nevertheless, CANET does not address OOD issues in other applications, such as molecular graphs. In this work, we propose a framework capable of generalizing to unseen environments characterized by differences not only in graph size but also in graph structure, node features, and edge features.
@@
Graph Data Augmentation. Beyond invariant graph learning, graph data augmentation aims to diversify the training distribution, thereby enhancing the OOD generalization of models. DropEdge (rong2019dropedge) introduces randomness by selectively removing edges, thus varying the training data’s structure. M-Mixup (wang2021mixup) enriches the dataset by interpolating diverse and irregular graphs within semantic space. $\mathcal{G}$-Mixup (han2022g) extends mixup to graph classification by interpolating across graph generators (graphons). Adversarial augmentation techniques, such as FLAG (kong2022robust), perturb node features via gradients, while AIA (sui2024unleashing) learns adversarial masks to probe environmental discrepancies. Despite these advances, methods that edit only the observed training graphs often have limited ability to explore truly novel environments and can distort label-determining patterns, yielding uncontrolled augmentations. Environment-aware frameworks (ligraph) partially address this by using environment metadata to linearly explore structures and features; however, they rely on high-quality, diverse environment labels that are costly and often infeasible to obtain. GRATIN (abbahaddou2025graph) takes a different route: it fits a Gaussian mixture on learned graph embeddings and samples synthetic representations—supported by a Rademacher-complexity analysis—to achieve efficient gains in generalization without editing raw graphs. In contrast, our work introduces a generation-based augmentation method that dispenses with environment labels and generated graphs end-to-end (structures, node features, and edge features), enabling controlled exploration beyond the training graph space while preserving stable, label-determining patterns.
3. Problem Formulation
@@
Notations. We represent a graph with $n$ nodes as $G=(\boldsymbol{A},\boldsymbol{X},\boldsymbol{E})$, where $\boldsymbol{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, $\boldsymbol{X}\in\mathbb{R}^{n\times a}$ denotes $a$-dimensional node features and $\boldsymbol{E}\in\mathbb{R}^{n\times n\times b}$ encodes $b$-dimensional edge features. Without the loss of generality, we focus on the graph classification task where each graph $G$ is associated with a label $Y\in\mathcal{Y}$, determined by a predefined labelling rule $\mathcal{G}\rightarrow\mathcal{Y}$. Following invariant learning (ahuja2021invariance; chen2022learning), we denote the graph dataset as $\mathcal{D}=\{(G_{i}^{e},Y_{i}^{e})\}_{e\in\mathcal{E}_{\mathrm{all}}}$, where $(G_{i}^{e},Y_{i}^{e})\sim P_{e}(G,Y)$ is an i.i.d. draw in the environment $e$ sampled from all possible environments $\mathcal{E}_{\mathrm{all}}$. The complete dataset can be partitioned into a training set $\mathcal{D}_{\mathrm{tr}}=\{(G_{i}^{e},Y_{i}^{e})\}_{e\in\mathcal{E}_{\mathrm{tr}}}$ and a test set $\mathcal{D}_{\mathrm{te}}=\{(G_{i}^{e},Y_{i}^{e})\}_{e\in\mathcal{E}_{\mathrm{te}}}$, where $\mathcal{E}_{\mathrm{tr}}$ and $\mathcal{E}_{\mathrm{te}}$ index the training and testing environments, respectively. In practice, environment information may not be explicitly given, and we further denote the training distribution as $P_{\mathrm{tr}}(G,Y)$ and the testing distribution as $P_{\mathrm{te}}(G,Y)$.
3.1. Graph Covariate Shift
We study OOD graph classification under covariate shift: the label mechanism is stable while the input distribution shifts, i.e.,
$P_{\mathrm{tr}}(Y\mid G)=P_{\mathrm{te}}(Y\mid G)\quad\text{but}\quad P_{\mathrm{tr}}(G)\neq P_{\mathrm{te}}(G).$
This mismatch arises from limited training coverage and changing test conditions. Prior work (ligraph) distinguishes: (i) feature shift $P_{\mathrm{tr}}(\mathbf{X})\neq P_{\mathrm{te}}(\mathbf{X})$ with $P_{\mathrm{tr}}(\mathbf{A},\mathbf{E})=P_{\mathrm{te}}(\mathbf{A},\mathbf{E})$ (e.g., GOOD-CMNIST (gui2022good), where color defines environments); (ii) structural shift $P_{\mathrm{tr}}(\mathbf{A},\mathbf{E})\neq P_{\mathrm{te}}(\mathbf{A},\mathbf{E})$; and (iii) joint shift $P_{\mathrm{tr}}(\mathbf{A},\mathbf{X},\mathbf{E})\neq P_{\mathrm{te}}(\mathbf{A},\mathbf{X},\mathbf{E})$ (e.g., GOOD-Motif size/basis changes (gui2022good), or scaffold splits in GOOD-HIV (gui2022good)). All three forms challenge OOD generalization; our framework is designed to address them.
3.2. Graph Classification under Covariate Shift
@@
With only observing the training set $\mathcal{D}_{\mathrm{tr}}$ sampled from the training distribution $P_{\mathrm{tr}}$ in training environments $\mathcal{E}_{\mathrm{tr}}$, our generalization objective under graph covariate shift is to train an optimal graph classifier $f:\mathcal{G}\rightarrow\mathcal{Y}$ that performs well across any possible environments $\mathcal{E}_{\mathrm{all}}\supseteq\mathcal{E}_{\mathrm{tr}}$. We formulate this goal as the following minimization problem:
(1) $\underset{f}{\min}\ \mathbb{E}_{e\in\mathcal{E}_{\mathrm{all}}}\mathbb{E}_{(G^{e},Y^{e})\sim P_{e}(G,Y)}[\ell(f(G^{e}),Y^{e})],$
where $\ell(\cdot,\cdot)$ denotes the loss function for graph classification and the expectation is with respect to graphs under all possible environments. However in practice, the training environments $\mathcal{E}_{\mathrm{tr}}$ may not cover all environments, causing degraded classification performance when applying the learned classifier in unseen test environments. This covariate shift calls for an effective manner to sufficiently explore unseen data distribution or environments during model training. We summarize and discuss the various types of graph covariate shifts in detail in Appendix.
3.3. Issues in Environmental Augmentation
To augment the training distribution for mitigating graph covariate shift, existing solutions often approach Eq. (1) by separating and augmenting the environments:
(2) ${\min_{f}}\ \mathbb{E}_{e\in\{\mathcal{E}_{\mathrm{tr}}\cup\mathcal{E}_{\mathrm{aug}}\}}\mathbb{E}_{(G^{e},Y^{e})\sim P_{e}(G,Y)}[\ell(f(G^{e}),Y^{e})],$
where the augmented environments $\mathcal{E}_{\mathrm{aug}}$ are obtained based on either interpolating explicitly given environmental labels (ligraph) or perturbing implicitly separated environmental components (sui2024unleashing; wu2022discovering; miao2022interpretable; chen2022learning). Obtaining accurate environmental labels and components itself could be high-cost and nontrivial tasks. Furthermore, without explicit environmental information and additional assumptions, separating environmental components could be inherently unfeasible (chen2024does), which limit the practicability of such strategy. Additionally, the subgraph perturbations based augmentation is mainly operated by edge dropping (sui2024unleashing; rong2019dropedge) and mixup (han2022g), which is confined to existing subgraphs in training data and could cause invalid samples (e.g., generating molecules that are chemically invalid).
4. Score-based Conditional Out-of-Distribution Diffusion Augmentation (CODA)
In this section we present CODA, a score-based augmentation framework that generates graph samples exploring new environments while preserving label-determining structure. We first introduce a distribution-level view of augmentation, then describe the forward and reverse diffusion processes, and finally formalize the working principles of CODA.
Our New Perspective: Distributional Augmentation with OOD Control. To overcome the limitations of existing strategies, this work formulates the augmentation problem as a generation-based graph OOD augmentation strategy, which directly models and augments the training distribution, without explicitly requiring the knowledge or separation of environmental information. Specifically, we target on synthesizing an augmented training distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$, which is combined with the original training distribution to obtain the classifier, stated as:
(3) $\underset{f}{\min}\ \mathbb{E}_{(G,Y)\sim\{P_{\mathrm{tr}}(G,Y)\cup\tilde{P}_{\mathrm{tr}}(G,Y)\}}[\ell(f(G),Y)].$
The augmented distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$ can be implemented in multiple ways, but it needs to satisfy two principles: (i) $\tilde{P}_{\mathrm{tr}}(G,Y)$ should deviate from $P_{\mathrm{tr}}(G,Y)$ in a controlled manner for exploration, and (ii) the explored graphs in $\tilde{P}_{\mathrm{tr}}(G,Y)$ should preserve the stable patterns of graphs in $P_{\mathrm{tr}}(G,Y)$. However, current graph generation models (jo2022score; martinkus2022spectre; vignac2022digress) cannot directly generate graphs that meet these two criteria. To address this, we propose a novel score-based conditional generative model that captures the augmented distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$ while adhering to both principles.